The max wavelength of radiation that can produce photoelectric effect in a certain metal is 200 nm. The max kinetic energy acquired by electron due to radiation of wavelength 100 nm will be? 1. 12.4ev2. 6.2 ev

Grade:12th pass

5 Answers

Adarsh
6 years ago
The answer will be 12.4ev .. 
Energy of wavelength (200-100) will give kinetic energy to electrons ..
so, Max. K.E= 12400/1000 = 12.4ev
Arun
6 years ago
Dear Kanika
 
Kmax = 12375 [1/1000  -  1/2000]
 = 6.2 ev
 
Hope it helps, in case of any query please feel free to ask
 
Regards
Arun (askIITians forum expert)
kanika
6 years ago
By putting what values you got 12375 in this? Please explain......... I was not able to get this value.
Krishangi
5 years ago
12375 is the value of hc
as KE=h√-h√° 
=> KE= hc/1000 A° - hc/2000A°
KE = hc(1/1000- 1/2000)
KE = 12375(1/1000-1/2000)
KE=6.2 
as hc=12375
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Kmax = 12375 [1/1000 – 1/2000]
 = 6.2 eV
 
Thanks and regards,
Kushagra